i need a solver for when the equation is not = 0.
Answer by stanbon(75887)

i need a solver for when the equation is not = 0.
---
If you have 2x^2 + 3x =4
you almost always subtract 4 from each side
to get 2x^2 + 3x -4 = 0 and then apply
the quadratic formula, or factor to find "x".
-----
If you leave the equation as 2x^2 + 3x = 4 you
can only solve it by guessing or by graphing.
=========================
Cheers,
Stan H.